### v2

# Quasi-invariant states with uniformly bounded cocycles

Abstract
We investigate the notion of quasi-invariant states introduced in [2, 3] from an analytic viewpoint. We give the structures of quasi-invariant states with uniformly bounded cocycles. As a consequence, we can apply a Theorem of Kovács and Szücs to get a conditional expectation on fixed points and another of $\o$rmer to get an invariant semifinite trace under extra assumptions.
1 Introduction
Let $(S,\mathcal{B})$ be a measurable space and $G$ a group of *-automorphisms acting on $S$ by measurable transformation
$s\in S\mapsto gs\in S,\quad g\in G.$
A measure $\mu$ on $(S,\mathcal{B})$ is called $G$-quasi-invariant if the measures $\mu\circ g$ and $\mu$ are absolutely continuous one with respect to the other.
Quasi-invariant measures are a much wider class than invariant, and that is why they constitute the natural environment for the theory of dynamical systems and ergodic theory. Therefore to develop a real quantum analogue of the classical theory of dynamical systems, requires developing a theory of quasi-invariant states on general operator algebras. Recently, in [2] the authors introduced the theory of quasi-invariant states under the action of a group of *-automorphisms on a von Neumann algebra (or a *-algebra). Moreover, in [3], they provide the first example of de Finetti Theorem for a quasi-exchangeable state (which is not exchangeable). The relationship between the group $G$ and modular automorphism group, invariant subalgebras, ergodicity, modular theory, and abelian subalgebras is studied in [7]. Finally, a martingale convergence theorem for strongly quasi-invariant states is given in [8].
In this note, we wish to clarify the notion $G$-quasi-invariant state under a natural boundedness assumption. The paper is structured as follows. First, we give the structures of quasi-invariant states with uniformly bounded cocycles. We further deal with a Theorem of Kovács and Szücs. Then, we extend the main result of Størmer in [19] to the strongly quasi-invariant states with uniformly bounded cocycles.
2 Quasi-invariant states
We briefly review some definitions and properties of quasi-invariant states, respectively, strongly quasi-invariant states with respect to a group of *-automorphisms on a von Neumann algebra (cf [2]).
Let $\mathcal{A}$ be a von Neumann algebra and $G$ be a group of normal *-automorphisms of $\mathcal{A}$ equipped with its usual strong-topology.
Definition 2.1
A faithful normal state $\varphi$ on $\mathcal{A}$ is said to be $G$-quasi-invariant if for all $g\in G$ there exists $x_{g}\in\mathcal{A}$ such that
$\varphi(g(a))=\varphi(x_{g}a),\quad a\in\mathcal{A}.$ (2.1)
We say that $\varphi$ is $G$-strongly quasi-invariant if it is $G$-quasi-invariant and the Radon-Nikodym derivatives $x_{g}$ are self-adjoint: $x_{g}=x_{g}^{*}$, $g\in G$.
It is proved in [2] the following properties:
  1. (i)
    When $\varphi$ is $G$-quasi-invariant, the map $g\mapsto x_{g}$ is a normalized multiplicative left $G$-1-cocycle satisfying
    $x_{e}=\hbox{\rm 1\kern-2.8ptl},\quad x_{g_{2}g_{1}}=x_{g_{1}}g_{1}^{-1}(x_{g_{2}}),\quad g_{1},g_{2}\in G,$ (2.2)
    and $x_{g}$ is invertible with its inverse
    $x_{g}^{-1}=g^{-1}(x_{g^{-1}}).$ (2.3)
    Furthermore, for all $a\in\mathcal{A}$ and $g\in G$ it holds
    $\varphi(x_{g}a)=\varphi(ax_{g}^{*}).$ (2.4)
  2. (ii)
    If $\varphi$ is $G$-strongly quasi-invariant, it holds that $x_{g}$ is strictly positive. It also holds that $x_{g}$ commutes with $x_{h}$ for all $g,h\in G$. Hence the von Neumann algebra $\mathcal{C}$ generated by $\{x_{g}:g\in G\}$ is commutative. It can be also easily shown that $\mathcal{C}$ is a subalgebra of $\mathrm{Centr}(\varphi)$, the centralizer of $\varphi$, which is defined by
    $\mathrm{Centr}(\varphi):=\{x\in\mathcal{A}:\varphi(xy)=\varphi(yx)\text{ for all }y\in\mathcal{A}\}$ (2.5)
Remark 2.2
If $\mathcal{A}$ is a von Neumann algebra and $\varphi$ and $\psi$ are two states on $\mathcal{A}$. We say that $\varphi$ is $\psi$–absolutely continuous if $\psi(a^{*}a)=0$ implies to $\varphi(a^{*}a)=0$ for $a\in\mathcal{A}$. We say that $\varphi$ and $\psi$ are equivalent ($\varphi\sim\psi$) if they are absolutely continuous one with respect to the other. A state $\varphi$ is $G$-invariant (resp. $G$–quasi-inavriant) if $\varphi\circ g=\varphi$ (resp. $\varphi\circ g\sim\varphi$) for all $g\in G$. Indeed, if $\varphi$ is as in Definition 2.1, then by [[18], 5.20. Lemma, page 124] one has
$\varphi\circ g(a^{*}a)=\varphi(x_{g}a^{*}a)\leq||x_{g}||\varphi(a^{*}a),\quad\varphi(a^{*}a)\leq||x_{g}^{-1}||\varphi_{g}(a^{*}a),\quad\forall g\in G,\;\forall a\in\mathcal{A},$
which proves that $\varphi\circ g\sim\varphi$ for all $g\in G$.
3 Quasi-invariant states with uniformly bounded cocycles
We will make the following assumption
Assumption. Assume that there exists $\lambda>0$ such that
$||x_{g}||\leq\lambda,\quad\forall g\in G.$ (A)
[…]
Lemma 3.1
Under assumption (A), for any $g\in G$ and any positive operator $a\in\mathcal{A}$
$\frac{1}{\lambda}\varphi(a)\leq\varphi(x_{g}a)\leq\lambda\varphi(a)$
Proof. Let $g\in G$ and $a\in\mathcal{A}$ be a positive operator. Then by assumption (A) and Remark 2.2
$0\leq\varphi(g(a))=\varphi(x_{g}a)\leq||x_{g}||\varphi(a)\leq\lambda\varphi(a)$
Moreover, one has
$\varphi(a)$ $=$ $\varphi(g^{-1}(g(a))$
$=$ $\varphi(x_{g^{-1}}(g(a))$
$\leq$ $||x_{g^{-1}}||\varphi(g(a))$
which proves that
$\varphi(g(a))\geq\frac{1}{||x_{g^{-1}}||}\varphi(a)\geq\frac{1}{\lambda}\varphi(a).$
$\Box$
Our main result states that this assumption is equivalent to have an invariant normal faithful state for the action. We borrow some ideas of [1].
Theorem 3.2
If $\varphi$ is a normal $G$-quasi-invariant state on $\mathcal{A}$ with cocycle $x_{g}$’s satisfying (A), then there exist a normal faithful $G$-invariant state $\psi$ on $\mathcal{A}$ and a bounded operator $d\in\mathcal{A}$ such that for all $g\in G$
$\psi(a):=\varphi(da),\quad a\in\mathcal{A}.$
Conversely, if there exist a normal faithful $G$-invariant state $\psi$ on $\mathcal{A}$ and a bounded invertible operator $d\in\mathcal{A}$ such that for all $g\in G$
$\psi(a):=\varphi(da),\quad a\in\mathcal{A},$
then $\varphi$ is a normal $G$-quasi-invariant state on $\mathcal{A}$ with cocycle $x_{g}$’s satisfying (A).
Proof. On $\mathcal{A}$, define the linear map
$\Gamma_{g}(a):=x_{g^{-1}}g(a),\quad a\in\mathcal{A}.$
Since $g$ is normal and $x_{g^{-1}}$ is bounded, it follows that $\Gamma_{g}$ is $\sigma$-weakly continuous. Moreover, the maps $\Gamma_{g}$ satisfy the following properties:
  1. (i)
    For all $g,h\in G$ , $\Gamma_{g}(x_{h})=x_{hg^{-1}}$. This is exactly (2.2).
  2. (ii)
    For all $h,g\in G$, one has $\Gamma_{gh}=\Gamma_{h}\Gamma_{g}$.
    If $a\in\mathcal{A}$, computing with (2.2)
    $\Gamma_{gh}(a)$ $=$ $x_{(gh)^{-1}}\;gh(a)$
    $=$ $x_{h^{-1}g^{-1}}\;gh(a)$
    $=$ $x_{g^{-1}}g(x_{h^{-1}})g(h(a))$
    $=$ $x_{g^{-1}}g(\Gamma_{h}(a))$
    $=$ $\Gamma_{g}(\Gamma_{h}(a)).$
  3. (iii)
    For all $g\in G$, $\varphi\circ\Gamma_{g}=\varphi$.
    If $a\in\mathcal{A}$, easy computations give
    $\varphi(\Gamma_{g}(a))=\varphi(x_{g^{-1}}g(a))=\varphi(g^{-1}(g(a)))=\varphi(a).$
  4. (iv)
    For all $g\in G$ and $a,b\in\mathcal{A}$, $\Gamma_{g}(ab)=\Gamma_{g}(a)x_{g^{-1}}^{-1}\Gamma_{g}(b)$.
  5. (v)
    For all $g\in G$ and $a\in\mathcal{A}$
    $(\Gamma_{g}(a))^{*}=x_{g^{-1}}^{-1}\Gamma_{g}(a^{*})(x_{g^{-1}})^{*}.$
Let $K$ be the $\sigma$-weak closure of $\mbox{conv}_{g\in G}\{x_{g}\}$. Since the sequence $\{x_{g}\}_{g\in G}$ is uniformly bounded, $K$ is convex and compact with respect to the $\sigma$-weak topology.
Moreover, by (i), $\Gamma_{g}(K)\subset K$ for all $g\in G$.
Now define a continuous norm $\rho$ on $\mathcal{A}$ by $\rho(x)=\varphi(xx^{*})^{1/2}$. Let $a\neq b\in K$, one has thanks to (i), (iv), (v) and Lemma 3.1
$\rho(\Gamma_{g}(a)-\Gamma_{g}(b))^{2}$ $=$ $\varphi(\Gamma_{g}(a-b)(\Gamma_{g}(a-b))^{*})$
$=$ $\varphi(\Gamma_{g}(a-b)x_{g^{-1}}^{-1}\Gamma_{g}((a-b)^{*})(x_{g^{-1}})^{*})$
$=$ $\varphi(\Gamma_{g}((a-b)(a-b)^{*})(x_{g^{-1}})^{*})$
[…]
$\geq$ $\frac{1}{\lambda}\varphi((a-b)(a-b)^{*})>0.$
Therefore $G$ is non-contracting and by the Ryll-Nardzewski fixed point theorem [16] there exists $d\in K$ such that $\Gamma_{g}(d)=d$, i.e. $d=x_{g^{-1}}g(d)$ for all $g\in G$. Notice that for all $g\in G$, $\varphi(x_{g})=1$. Then for all $x\in K$
$\varphi(x)=1=\varphi(d).$
Since $\varphi$ is a faithful state, the operator $d$ is non zero. Define
$\psi(a):=\varphi(da),\quad a\in\mathcal{A}.$
Then for all $g\in G$
$\psi(g(a))=\varphi(dg(a))=\varphi\big{(}g(g^{-1}(d)a)\big{)}=\varphi\big{(}x_{g}g^{-1}(d)a\big{)}=\varphi(da)=\psi(a)$
which proves that $\psi$ is $G$-invariant. By Lemma 3.1 for any $y\in K$, we have $\frac{1}{\lambda}\varphi\leq\varphi.y\leq\lambda\varphi$. In particular $\psi$ is faithful as $\varphi$ is.
Conversely, let $d$ be a bounded invertible element of $\mathcal{A}$ such that
$\psi(a):=\varphi(da),\quad a\in\mathcal{A}$
is $G$-invariant. It follows that for all $a\in\mathcal{A}$, $\varphi(a)=\psi(d^{-1}a)$, and for all $g\in G$
$\varphi(g(a))$ $=$ $\psi(d^{-1}g(a))$
$=$ $\psi(g(g^{-1}(d^{-1})a))$
$=$ $\psi(g^{-1}(d^{-1})a)$
$=$ $\varphi(dg^{-1}(d^{-1})a).$
This proves that $\varphi$ is $G$-quasi-invariant state with cocycle given by
$x_{g}=dg^{-1}(d^{-1}).$ (3.1)
Since $d$ is bounded invertible operator, there exists $M>0$ such that $d^{-1}(d^{-1})^{*}\leq M$. Therefore for any $g\in G$
$g^{-1}(d^{-1}(d^{-1})^{*})\leq M\;\;\mbox{and}\;\;d^{-1}g^{-1}(d^{-1}(d^{-1})^{*})(d^{-1})^{*}\leq M^{2}.$
It follows that for any $g\in G$
$||x_{g}||^{2}=||x_{g}x_{g}^{*}||\leq M^{2}.$
Thus the family $\{x_{g}\}_{g\in G}$ satisfies (A). $\Box$
Now, let $\varphi$ be a normal faithful $G$-strongly quasi-invariant state on $\mathcal{A}$ with cocycle $x_{g}^{\prime}$s satisfying assumption (A). Since $x_{g}$ is a positive operator, one has
$x_{g}\leq\lambda,\quad\forall g\in G,$
which is equivalent to
$x_{g}^{-1}\geq\frac{1}{\lambda},\quad\forall g\in G.$ (3.2)
Notice that for all $g\in G$, $x_{g}^{-1}=g(x_{g^{-1}})$. Then from (3.2) it follows that
$x_{g^{-1}}\geq\frac{1}{\lambda},\quad\forall g\in G$
and
$x_{g}\geq\frac{1}{\lambda},\quad\forall g\in G.$
Therefore it it is clear that assumption (A) is equivalent to
$\frac{1}{\lambda}\leq x_{g}\leq\lambda,\quad\forall g\in G.$ (3.3)
As a consequence of Theorem 3.2, we prove the following
Theorem 3.3
$\varphi$ is a normal faithful $G$-strongly quasi-invariant state with cocycle $x_{g}$’s satisfying (A) if and only if there exist a normal faithful $G$-invariant state $\psi$ and a bounded positive invertible operator $d\in\mathcal{A}$ such that for all $g\in G$
$\psi(a):=\varphi(da),\quad a\in\mathcal{A}.$
Moreover, in this case for all $g\in G$, $[d,\;g(d)]=0$.
Proof. Since $\varphi$ is a normal faithful $G$–strongly quasi-invariant state with cocycle $x_{g}$’s satisfying assumption (A) then by Theorem 3.2, there exist a normal faithful $G$-invariant state $\psi$ and a bounded operator $d\in\mathcal{A}$ such that for all $g\in G$
$\psi(a):=\varphi(da),\quad a\in\mathcal{A}.$
Remember that from the proof of Theorem 3.2, $K$ is the $\sigma$-weak closure of $\mbox{conv}_{g\in G}\{x_{g}\}$ and the cocycle $x_{g}$’s satisfy (3.3), it follows that $d$ satisfies (3.3) and $d$ is a positive invertible operator in $\mathcal{A}$. Notice that in this case $x_{g}=dg^{-1}(d^{-1})=x_{g}^{*}$ which is equivalent to $[d,\;g(d)]=0$.
Conversely, let $\psi$ be a normal faithful $G$-invariant state
$\psi(a):=\varphi(da),\quad a\in\mathcal{A}$
for some normal state $\varphi$ on $\mathcal{A}$, where $d$ is a positive invertible operator in $\mathcal{A}$ such that for all $g\in G$, $[d,\;g(d)]=0$. Then it is clear that $\varphi$ is a normal faithful $G$–strongly quasi-invariant state on $\mathcal{A}$ with cocycle $x_{g}=dg^{-1}(d^{-1})=x_{g}^{*}$. Since $d$ is a positive invertible operator, it is clear that the cocycles $x_{g}^{\prime}$s satisfy (A). $\Box$
4 Unitary transformations associated to quasi–invariant states
When $\varphi$ is $G$-strongly quasi-invariant state on a von Neumann algebra $\mathcal{A}$ (for a given action) with cocycle $x_{g}$’s, it is proved in [2] that there exists a unique unitary representation $U$ of $G$ on $\mathcal{H}$ implementing $\pi$. We will show in this section that for $G$-quasi-invariant states, one can always spatially implement the action of $G$ (but we don’t know if the map $g\mapsto U_{g}$ is a representation).
Let $G$ be a group of normal *-automorphisms of a von Neumann algebra $\mathcal{A}$ and $\varphi$ be a normal $G$-quasi-invariant state on $\mathcal{A}$ with cocycle $x_{g}^{\prime}$s.
We will use Haagerup formalism in this section for the standard form of $\mathcal{A}$ as in section II of [22], this will avoid heavy formulas. We denote by $\mathcal{N}$ the crossed product $\mathcal{A}\rtimes\sigma^{\varphi}$ of $\mathcal{A}$ by the modular group $\sigma^{\varphi}$ associated with $\varphi$. We denote by $\theta$ the dual action of $\mathbb{R}$ on $\mathcal{N}$. Then $\theta$ satisfies $\theta_{s}(x)=x\;\;(s\in\mathbb{R})$ iff $x\in\mathcal{A}$. Let $\tau$ be the trace on $\mathcal{N}$ such that $\tau\circ\theta_{s}=e^{-s}\tau$ and denote by $L_{0}(\mathcal{N},\;\tau)$ the space of $\tau$-measurable elements $x\in\mathcal{N}$. Define for $p=1,2$
$L_{p}(\mathcal{A}):=\{x\in L_{0}(\mathcal{N},\;\tau)|\;\;\theta_{s}(x)=e^{-s/p}x,\;\;s\in\mathbb{R}\}.$
There is an identification of $\mathcal{A}_{*}$ with $L_{1}(\mathcal{A})$ assigning a density $D_{\gamma}\in L_{1}(\mathcal{A})$ to a normal linear form $\gamma\in\mathcal{A}_{*}$ (Proposition 15 [22]). The duality is defined through the use of the trace $\mbox{Tr}:L_{1}(\mathcal{A})\to\mathbb{C}$ with
$\langle D_{\gamma},\;a\rangle_{L_{1}(\mathcal{A}),\;\mathcal{A}}=\hbox{\rm{Tr}}(D_{\gamma}a)=\gamma(a).$
Since $\varphi$ is faithful state on $\mathcal{A}$, its density $D_{\varphi}\in L_{1}(\mathcal{A})$ satisfies supp$(D_{\varphi})=1$ (Lemma 2, [22]). Note that by duality $G$ acts on $L_{1}(\mathcal{A})$ and we have
$\langle g^{*}D_{\varphi},\;a\rangle_{L_{1}(\mathcal{A}),\,\mathcal{A}}=\varphi(g(a)).$
The space $L_{2}(\mathcal{A})$ is naturally a Hilbert space with scalar product $\langle A,B\rangle=\hbox{\rm{Tr}}(A^{*}B)$. Moreover, by Theorem 36 in [22], the GNS representation of $(\mathcal{A},\;\varphi)$ is the triple $(\mathcal{A}\subset\mathcal{N},\;L_{2}(\mathcal{A}),\;D_{\varphi}^{1/2})$.
Lemma 4.1
There exists a positive invertible operator $a_{g}$ in $\mathcal{A}$ such that
$D_{\varphi\circ g}=D_{\varphi}^{1/2}a_{g}^{2}D_{\varphi}^{1/2}=x_{g}^{*}D_{\varphi}.$ (4.1)
Proof. Since $\varphi$ is normal $G$-quasi-invariant and for all $a\in\mathcal{A}$ and $g\in G$,
$\varphi(x_{g}a)=\varphi(ax_{g}^{*})$, it follows that
$g^{*}D_{\varphi}:=D_{\varphi\circ g}=x_{g}^{*}D_{\varphi}=D_{\varphi}x_{g}.$ (4.2)
On the other hand, for any positive operator $a\in\mathcal{A}$ and any $g\in G$, one has
$\varphi(g(a))=\hbox{\rm{Tr}}(D_{\varphi\circ g}\,a)=\varphi(x_{g}a)\leq||x_{g}||\varphi(a)=||x_{g}||\,\hbox{\rm{Tr}}(D_{\varphi}^{1/2}aD_{\varphi}^{1/2}).$
By Proposition 4 [22], we have $0\leq D_{\varphi\circ g}\leq\|x_{g}\|D_{\varphi}$. As $D_{\varphi}$ has full support, there exists a positive element $a_{g}$ of $\mathcal{N}$ such that
$D_{\varphi\circ\,g}=D_{\varphi}^{1/2}a_{g}^{2}D_{\varphi}^{1/2}.$
Considering the dual action yields that $\theta_{s}(a_{g})=a_{g}$ for all $s\in\mathbb{R}$ and $a_{g}$ is a positive operator in $\mathcal{A}$. Now, notice that for any positive operator $a$ in $\mathcal{A}$
$\varphi(a)=\varphi\circ g(x_{g}^{-1}a)\leq||x_{g}^{-1}||\,\varphi\circ g(a).$
Hence one gets
$D_{\varphi\circ g}=D_{\varphi}^{1/2}a_{g}^{2}D_{\varphi}^{1/2}\geq\alpha D_{\varphi},$
where $\alpha=||x_{g}^{-1}||^{-1}$. This proves that $a_{g}^{2}\geq\alpha$ and $a_{g}$ is invertible. $\Box$
Actually, the equation $x_{g}^{*}D_{\varphi}=D_{\varphi}x_{g}$ implies that $x_{g}$ sits in the domain of $\sigma^{\varphi}_{-i}$ and $\sigma^{\varphi}_{-i}(x_{g})=x_{g}^{*}$ (see Theorem 3.25 in [21]). In particular, $x_{g}$ sits in the domain of $\sigma^{\varphi}_{-i/2}$ and one can check that
$a_{g}^{2}=\sigma^{\varphi}_{-i/2}(x_{g}).$ (4.3)
Now we prove the following
Theorem 4.2
For all $g\in G$, the map
$U_{g}:$ $\,\mathcal{A}\,D_{\varphi}^{1/2}\subset L_{2}(\mathcal{A})\longrightarrow L_{2}(\mathcal{A})$
$xD_{\varphi}^{1/2}\longmapsto g^{-1}(x)D_{\varphi}^{1/2}a_{g}$
extends to a unitary transformation on $L_{2}(\mathcal{A})$ (still denoted by $U_{g}$). Moreover for any $x\in\mathcal{A}$
$U_{g}^{*}xU_{g}=g(x).$
Proof. For all $x,y\in\mathcal{A}$, one has
$\langle U_{g}(xD_{\varphi}^{1/2}),\,U_{g}(yD_{\varphi}^{1/2})\rangle_{L_{2}(\mathcal{A}),\,L_{2}(\mathcal{A})}$ $=$ $\hbox{\rm{Tr}}(a_{g}D_{\varphi}^{1/2}g^{-1}(x)^{*}g^{-1}(y)D_{\varphi}^{1/2}a_{g})$
$=$ $\hbox{\rm{Tr}}(D_{\varphi}^{1/2}a_{g}^{2}D_{\varphi}^{1/2}g^{-1}(x)^{*}g^{-1}(y))$
$=$ $\varphi(x^{*}y)$
$=$ $\langle xD_{\varphi}^{1/2},\,yD_{\varphi}^{1/2}\rangle_{L_{2}(\mathcal{A}),\,L_{2}(\mathcal{A})}.$
Since $\mathcal{A}\,D_{\varphi}^{1/2}$ is dense in $L_{2}(\mathcal{A})$, $U_{g}$ can be extended to an isometry on $L_{2}(\mathcal{A})$. Now, let $h\in L_{2}(\mathcal{A})\cap\mbox{Ran}(U_{g})$. Then, for all $a\in\mathcal{A}$, $\hbox{\rm{Tr}}(a_{g}h^{*}g^{-1}D_{\varphi}^{1/2})=0$. Notice that $\mathcal{A}\,D_{\varphi}^{1/2}$ is dense in $L_{2}(\mathcal{A})$ and $a_{g}$ is invertible. Therefore, one gets $a_{g}h^{*}=0$ and $h=0$. Hence, $U_{g}$ is a unitary operator on $L_{2}(\mathcal{A})$.
Now, let $x,y,z\in\mathcal{A}$. Then, one has
$\langle U_{g}^{*}xU_{g}(yD_{\varphi}^{1/2}),\,zD_{\varphi}^{1/2})\rangle_{L_{2}(\mathcal{A}),\,L_{2}(\mathcal{A})}$ $=$ $\langle xg^{-1}(y)D_{\varphi}^{1/2}a_{g},\;g^{-1}(z)D_{\varphi}^{1/2}a_{g}\rangle_{L_{2}(\mathcal{A}),\,L_{2}(\mathcal{A})}$
$=$ $\hbox{\rm{Tr}}(a_{g}D_{\varphi}^{1/2}(g^{-1}(y))^{*}x^{*}g^{-1}(z)D_{\varphi}^{1/2}a_{g})$
$=$ $\varphi(y^{*}g(x)^{*}z))$
$=$ $\langle g(x)yD_{\varphi}^{1/2},\;zD_{\varphi}^{1/2})\rangle_{L_{2}(\mathcal{A}),\,L_{2}(\mathcal{A})},$
which proves that $U_{g}^{*}xU_{g}=g(x).$Box
In the setting of Theorem 3.2, we have that there exists an invertible $a\in\mathcal{A}$ such that $D_{\psi}=D_{\varphi}^{1/2}a^{2}D_{\varphi}^{1/2}$. Thus there is a partial unitary $v$ such that $D_{\psi}^{a/2}=vaD_{\varphi}^{1/2}$. Since both operators have full support, $v$ is a unitary and we deduce that there exists an invertible $\gamma\in\mathcal{A}$ such that $D_{\psi}^{1/2}=\gamma D_{\varphi}^{1/2}$. Thus we get $D_{\psi}=dD_{\varphi}=\gamma D_{\varphi}\gamma^{*}$. As above, we get that $\gamma^{*}$ sits in the domain of $\sigma^{\varphi}_{-i}$ and $\sigma^{\varphi}_{-i}(\gamma^{*})=\gamma d$. We conclude that $d=\gamma^{-1}\sigma^{\varphi}_{-i}(\gamma^{*})$.
In the same way, let $\gamma_{g}=\gamma^{-1}g^{-1}(\gamma)$ for $g\in G$. then one can also check that $\gamma_{g}^{*}$ sits in the domain of $\sigma^{\varphi}_{-i}$ and that $x_{g}=\gamma_{g}^{-1}\sigma^{\varphi}_{-i}(\gamma_{g}^{*})$.
Proposition 4.3
If $\varphi$ is $G$-strongly quasi-invariant then $g\mapsto U_{g}$ is a unitary representation of the group $G$.
Proof. Notice that if $\varphi$ is $G$-strongly quasi-invariant, it is showed in [2] that the cocycles $x_{g},\;g\in G$, are in the centralizer of $\varphi$. Therefore by (4.3) it follows that $a^{2}_{g}=x_{g}=x_{g}^{*}$ (i.e. $a_{g}=x_{g}^{1/2}$). Moreover from (4.2), one has
$a_{g}D_{\varphi}^{1/2}=D_{\varphi}^{1/2}a_{g}$
and hence for all $x\in\mathcal{A}$, $g\in G$
$U_{g}(xD_{\varphi}^{1/2})=g^{-1}(x)x_{g}^{1/2}D_{\varphi}^{1/2}.$
By using the cocycle property it is easy to check that $U_{g}U_{h}=U_{gh}$ for all $g,h\in G$ and thus $g\mapsto U_{g}$ is a representation of $G$, which satisfies $U_{g}^{*}=U_{g}^{-1}$. $\Box$
5 A Theorem of Kovács and Szücs for quasi–invariant states with uniformly bounded cocycles
By Theorem 3.2, if $\varphi$ is a normal faithful $G$-quasi–invariant state on a von Neumann algebra $\mathcal{A}$ with cocycle satisfying assumption (A), then there exist a normal faithful $G$-invariant state $\psi$ on $\mathcal{A}$ and a bounded operator $d\in\mathcal{A}$ such that for all $g\in G$
$\psi(a):=\varphi(da),\quad a\in\mathcal{A}.$
Since $\psi$ is a normal faithful $G$-invariant state on $\mathcal{A}$, it follows from [9, Theorem 1] that there exists a normal faithful $G$-invariant Umegaki conditional expectation $\Phi:\mathcal{A}\rightarrow\mathcal{B}$, where $\mathcal{B}$ is a sub-von Neumann algebra of $\mathcal{A}$, such that
$\psi(.)=(\psi|\mathcal{B})\circ\Phi(.)$
If moreover $d$ is invertible, one gets
$\varphi=(\varphi(d.)|\mathcal{B})\circ\Phi(d^{-1}.)$ (5.1)
Now, if $\varphi$ is a $G$-strongly quasi-invariant state and if $g\mapsto U_{g}$ is the unitary representation of $G$ defined by Theorem 4.2, we will show that the conditional expectation $\Phi$ in (5.1) can also be characterized by $U_{g},\;g\in G$ Define
$u_{g}(a):=U^{*}_{g}aU_{g}=g(a)\ ;\quad\forall\,g\in G\ ;\quad\forall\,a\in\mathcal{A}$ (5.2)
Let $E_{0}$ be the orthogonal projection onto the subspace of all $\xi\in L_{2}(\mathcal{A})$ invariant under all $U_{g}$. We denote by $\mathcal{B}$ the set of fixed points of $\mathcal{A}$ under the action of the automorphisms $u_{g}$, $g\in G$. Let $F_{0}=[\mathcal{B}^{\prime}E_{0}]$ be the orthogonal projection onto $\overline{\mbox{span}(\mathcal{B}^{\prime}E_{0}L_{2}(\mathcal{A}))}$. Finally, let $\mathcal{M}$ be a Godement mean over $G$. The following result follows from a theorem of Kovács and Szücs (see [14] or [1] and [9, Theorem 1]).
Theorem 5.1
Assume that $\varphi$ is a faithful normal $G$-strongly quasi-invariant state with uniformly bounded cocycle (i.e. cocycle satisfying (A)). Then
  1. (1)
    There exists a unique normal faithful $G$-invariant Umegaki conditional expectation $\Phi:\mathcal{A}\rightarrow\mathcal{B}$.
    Moreover, $\Phi(b)$ can be defined (equivalently) as
    1. (i)
      the unique element of $\mathcal{A}$ such that Φ(b)E0