### v3

# Quasi-invariant states with uniformly bounded cocycles

Abstract
We investigate the notion of quasi-invariant states introduced in [2] from an analytic viewpoint. We give the structures of quasi-invariant states with uniformly bounded cocycles. As a consequence, we can apply a Theorem of Kovács and Szücs to get a conditional expectation on fixed points and another of $\o$rmer to get an invariant semifinite trace under extra assumptions.
1 Introduction
Let $(S,\mathcal{B})$ be a measurable space and $G$ a group of *-automorphisms acting on $S$ by measurable transformation
$s\in S\mapsto gs\in S,\quad g\in G.$
A measure $\mu$ on $(S,\mathcal{B})$ is called $G$-quasi-invariant if the measures $\mu\circ g$ and $\mu$ are absolutely continuous one with respect to the other.
Quasi-invariant measures are a much wider class than invariant, and that is why they constitute the natural environment for the theory of dynamical systems and ergodic theory. Therefore to develop a real quantum analogue of the classical theory of dynamical systems, requires developing a theory of quasi-invariant states on general operator algebras. Recently, in [2] the authors introduced the theory of quasi-invariant states under the action of a group of *-automorphisms on a von Neumann algebra (or a *-algebra). The relationship between the group $G$ and modular automorphism group, invariant subalgebras, ergodicity, modular theory, and abelian subalgebras is studied in [6]. Moreover, a martingale convergence theorem for strongly quasi-invariant states is given in [7]. Finally, in [15], quasi-invariant actions are used in the setting of discrete quantum groups. In particular, the main technical results in [15] is the computation of the Radon-Nikodym cocycle for the action of the dual of $SUq(2)$ on its Martin boundary with respect to the canonical harmonic state defined by a quantum random walk.
In this note, we wish to clarify the notion of $G$-quasi-invariant state under a natural boundedness assumption. The paper is structured as follows. First, we give the structure of quasi-invariant states with uniformly bounded cocycle. Moreover, for $G$-quasi-invariant states, we show that one can always spatially implement the action of $G$. We further deal with a Theorem of Kovács and Szücs. Then, we extend the main result of Størmer in [20] to strongly quasi-invariant states with uniformly bounded cocycles.
2 Quasi-invariant states
We briefly review some definitions and properties of quasi-invariant states, respectively, strongly quasi-invariant states with respect to a group of *-automorphisms on a von Neumann algebra (cf [2]).
Let $\mathcal{A}$ be a von Neumann algebra and $G$ be a group of normal *-automorphisms of $\mathcal{A}$ equipped with its usual strong-topology.
Definition 2.1
A faithful normal state $\varphi$ on $\mathcal{A}$ is said to be $G$-quasi-invariant if for all $g\in G$ there exists $x_{g}\in\mathcal{A}$ such that
$\varphi(g(a))=\varphi(x_{g}a),\quad a\in\mathcal{A}.$ (2.1)
We say that $\varphi$ is $G$-strongly quasi-invariant if it is $G$-quasi-invariant and the Radon-Nikodym derivatives $x_{g}$ are self-adjoint: $x_{g}=x_{g}^{*}$, $g\in G$.
It is proved in [2] the following properties:
  1. (i)
    When $\varphi$ is $G$-quasi-invariant, the map $g\mapsto x_{g}$ is a normalized multiplicative left $G$-1-cocycle satisfying
    $x_{e}=\hbox{\rm 1\kern-2.8ptl},\quad x_{g_{2}g_{1}}=x_{g_{1}}g_{1}^{-1}(x_{g_{2}}),\quad g_{1},g_{2}\in G,$ (2.2)
    and $x_{g}$ is invertible with its inverse
    $x_{g}^{-1}=g^{-1}(x_{g^{-1}}).$ (2.3)
    Furthermore, for all $a\in\mathcal{A}$ and $g\in G$ it holds
    $\varphi(x_{g}a)=\varphi(ax_{g}^{*}).$ (2.4)
  2. (ii)
    If $\varphi$ is $G$-strongly quasi-invariant, it holds that $x_{g}$ is positive invertible. It also holds that $x_{g}$ commutes with $x_{h}$ for all $g,h\in G$. Hence the von Neumann algebra $\mathcal{C}$ generated by $\{x_{g}:g\in G\}$ is commutative. It can be also easily shown that $\mathcal{C}$ is a subalgebra of $\mathrm{Centr}(\varphi)$, the centralizer of $\varphi$, which is defined by
    $\mathrm{Centr}(\varphi):=\{x\in\mathcal{A}:\varphi(xy)=\varphi(yx)\text{ for all }y\in\mathcal{A}\}.$ (2.5)
Now we introduce the following lemma which plays a crucial rule later.
Lemma 2.2 ([19], 5.20. Lemma)
Let $\varphi$ be a positive form on a $C^{*}$-algebra $\mathcal{A}$ and $a\in\mathcal{A}$. If the linear form $L_{a}\varphi$ defined on $\mathcal{A}$ by $L_{a}\varphi(x)=\varphi(ax)$ is positive, then
$L_{a}\varphi(x)\leq||a||\varphi(x),$
for all positive element $x\in\mathcal{A}$.
Remark 2.3
If $\mathcal{A}$ is a von Neumann algebra and $\varphi$, $\psi$ are two states on $\mathcal{A}$, we say that $\varphi$ is $\psi$–absolutely continuous if $\psi(a^{*}a)=0$ implies to $\varphi(a^{*}a)=0$ for $a\in\mathcal{A}$. We say that $\varphi$ and $\psi$ are equivalent ($\varphi\sim\psi$) if they are absolutely continuous one with respect to the other. A state $\varphi$ is $G$-invariant (resp. $G$–quasi-inavriant) if $\varphi\circ g=\varphi$ (resp. $\varphi\circ g\sim\varphi$) for all $g\in G$. Indeed, if $\varphi$ is as in Definition 2.1, then by Lemma 2.2 one has
$\varphi\circ g(a^{*}a)=\varphi(x_{g}a^{*}a)\leq||x_{g}||\varphi(a^{*}a),\quad\varphi(a^{*}a)\leq||x_{g}^{-1}||\varphi\circ g(a^{*}a),\quad\forall g\in G,\;\forall a\in\mathcal{A},$
which proves that $\varphi\circ g\sim\varphi$ for all $g\in G$.
3 Quasi-invariant states with uniformly bounded cocycles
We will make the following assumption.
Assumption. Assume that there exists $\lambda>0$ such that
$||x_{g}||\leq\lambda,\quad\forall g\in G.$ (A)
[…]
Lemma 3.1
Under assumption (A), for any $g\in G$ and any positive operator $a\in\mathcal{A}$
$\frac{1}{\lambda}\varphi(a)\leq\varphi(x_{g}a)\leq\lambda\varphi(a)\quad\textrm{and}\quad\frac{1}{\lambda}\varphi(a)\leq\varphi(a(x_{g}^{-1})^{*})\leq\lambda\varphi(a).$
Proof. Let $g\in G$ and $a\in\mathcal{A}$ be a positive operator. Then by assumption (A) and Remark 2.3
$0\leq\varphi(g(a))=\varphi(x_{g}a)\leq||x_{g}||\varphi(a)\leq\lambda\varphi(a).$
Moreover, one has
$\varphi(a)$ $=$ $\varphi(g^{-1}(g(a))$
$=$ $\varphi(x_{g^{-1}}(g(a))$
$\leq$ $||x_{g^{-1}}||\varphi(g(a))$
which proves that
$\varphi(g(a))\geq\frac{1}{||x_{g^{-1}}||}\varphi(a)\geq\frac{1}{\lambda}\varphi(a).$
Since $\varphi_{g}=\varphi\circ g:a\mapsto\varphi(x_{g}a)$ is a state, the linear form $a\mapsto\varphi(a(x_{g}^{-1})^{*})=\varphi_{g}(x_{g}^{-1}a(x_{g}^{-1})^{*})$ is positive. Then, for all $a\in\mathcal{A}$, we have
$\varphi(a(x_{g}^{-1})^{*})=\overline{\varphi(a^{*}(x_{g}^{-1})^{*})}=\overline{\varphi((x_{g}^{-1}a)^{*})}=\varphi(x_{g}^{-1}a)=L_{x_{g}^{-1}}\varphi(a)$
thanks to the positivity of the above-mentioned linear form and $\varphi$, and we can use Lemma 2.2 again as $L_{x_{g}^{-1}}\varphi\geq 0$. Similarly $\varphi=L_{x_{g}}(L_{x_{g}^{-1}}\varphi)\geq 0$ gives the last inequality. $\Box$
Our main result states that this assumption is equivalent to have an invariant normal faithful state for the action. We borrow some ideas of [1].
Theorem 3.2
If $\varphi$ is a normal $G$-quasi-invariant state on $\mathcal{A}$ with cocycle $x_{g}$’s satisfying (A), then there exist a normal faithful $G$-invariant state $\psi$ on $\mathcal{A}$ and a bounded operator $d\in\mathcal{A}$ such that
$\psi(a)=\varphi(da),\quad a\in\mathcal{A}.$
Conversely, if there exist a normal faithful $G$-invariant state $\psi$ on $\mathcal{A}$ and a bounded invertible operator $d\in\mathcal{A}$ such that for all $g\in G$
$\psi(a)=\varphi(da),\quad a\in\mathcal{A},$
then $\varphi$ is a normal $G$-quasi-invariant state on $\mathcal{A}$ with cocycle $x_{g}$’s satisfying (A).
Proof. On $\mathcal{A}$, define the linear map
$\Gamma_{g}(a):=x_{g^{-1}}g(a),\quad a\in\mathcal{A}.$
Since $g$ is normal and $x_{g^{-1}}$ is bounded, it follows that $\Gamma_{g}$ is $\sigma$-weakly continuous. Moreover, the maps $\Gamma_{g}$ satisfy the following properties:
  1. (i)
    For all $g,h\in G$ , $\Gamma_{g}(x_{h})=x_{hg^{-1}}$. This is exactly (2.2).
  2. (ii)
    For all $h,g\in G$, one has $\Gamma_{gh}=\Gamma_{g}\Gamma_{h}$.
    If $a\in\mathcal{A}$, computing with (2.2)
    $\Gamma_{gh}(a)$ $=$ $x_{(gh)^{-1}}\;gh(a)$
    $=$ $x_{h^{-1}g^{-1}}\;gh(a)$
    $=$ $x_{g^{-1}}g(x_{h^{-1}})g(h(a))$
    $=$ $x_{g^{-1}}g(\Gamma_{h}(a))$
    $=$ $\Gamma_{g}(\Gamma_{h}(a)).$
  3. (iii)
    For all $g\in G$, $\varphi\circ\Gamma_{g}=\varphi$.
    If $a\in\mathcal{A}$, easy computations give
    $\varphi(\Gamma_{g}(a))=\varphi(x_{g^{-1}}g(a))=\varphi(g^{-1}(g(a)))=\varphi(a).$
  4. (iv)
    For all $g\in G$ and $a,b\in\mathcal{A}$, $\Gamma_{g}(ab)=\Gamma_{g}(a)(x_{g^{-1}})^{-1}\Gamma_{g}(b)$.
  5. (v)
    For all $g\in G$ and $a\in\mathcal{A}$
    $(\Gamma_{g}(a))^{*}=(x_{g^{-1}})^{-1}\Gamma_{g}(a^{*})(x_{g^{-1}})^{*}.$
Let $K$ be the $\sigma$-weak closure of $\mbox{conv}\{x_{g}\}_{g\in G}$. Since the set $\{x_{g}\}_{g\in G}$ is uniformly bounded, $K$ is convex and compact with respect to the $\sigma$-weak topology.
Moreover, by (i), $\Gamma_{g}(K)\subset K$ for all $g\in G$.
Now define a continuous norm $\rho$ on $\mathcal{A}$ by $\rho(x)=\varphi(xx^{*})^{1/2}$. Let $a\neq b\in K$, thanks to the second estimate in Lemma 3.1 applied to $\Gamma_{g}(a-b)(\Gamma_{g}(a-b))^{*}\geq 0$ for $g^{-1}$:
$\rho(\Gamma_{g}(a)-\Gamma_{g}(b))^{2}\geq\frac{1}{\lambda}\varphi(\Gamma_{g}(a-b)(\Gamma_{g}(a-b))^{*}((x_{g^{-1}})^{-1})^{*}).$
Using the algebraic identities (v), (iv) and (iii)
$\lambda\rho(\Gamma_{g}(a)-\Gamma_{g}(b))^{2}$$\geq$ $\varphi(\Gamma_{g}(a-b)(x_{g^{-1}})^{-1}\Gamma_{g}((a-b)^{*}))$
$=$ $\varphi(\Gamma_{g}((a-b)(a-b)^{*}))$
$=$ $\varphi((a-b)(a-b)^{*})>0.$
Therefore $\Gamma_{G}$ is non-contracting and by the Ryll-Nardzewski fixed point theorem [17] there exists $d\in K$ such that $\Gamma_{g}(d)=d$, i.e. $d=x_{g^{-1}}g(d)$ for all $g\in G$. Notice that for all $g\in G$, $\varphi(x_{g})=1$. Then for all $x\in K$
$\varphi(x)=1=\varphi(d).$
It follows that the operator $d$ is non zero. Define
$\psi(a):=\varphi(da),\quad a\in\mathcal{A}.$
Then for all $g\in G$
$\psi(g(a))=\varphi(dg(a))=\varphi\big{(}g(g^{-1}(d)a)\big{)}=\varphi\big{(}x_{g}g^{-1}(d)a\big{)}=\varphi(da)=\psi(a)$
which proves that $\psi$ is $G$-invariant. By Lemma 3.1 for any $y\in K$, we have $\frac{1}{\lambda}\varphi\leq\varphi.y\leq\lambda\varphi$. In particular $\psi$ is faithful as $\varphi$ is.
Conversely, let $d$ be a bounded invertible element of $\mathcal{A}$ such that
$\psi(a):=\varphi(da),\quad a\in\mathcal{A}$
is $G$-invariant. It follows that for all $a\in\mathcal{A}$, $\varphi(a)=\psi(d^{-1}a)$, and for all $g\in G$
$\varphi(g(a))$ $=$ $\psi(d^{-1}g(a))$
$=$ $\psi(g(g^{-1}(d^{-1})a))$
$=$ $\psi(g^{-1}(d^{-1})a)$
$=$ $\varphi(dg^{-1}(d^{-1})a).$
This proves that $\varphi$ is $G$-quasi-invariant state with cocycle
$x_{g}=dg^{-1}(d^{-1}).$ (3.1)
Since $d$ is a bounded invertible operator, then for any $g\in G$
$g^{-1}(d^{-1}(d^{-1})^{*})\leq||d^{-1}||^{2}\;\;\mbox{and}\;\;dg^{-1}(d^{-1}(d^{-1})^{*})d^{*}\leq||d||^{2}||d^{-1}||^{2}.$
It follows that for any $g\in G$
$||x_{g}||^{2}=||x_{g}x_{g}^{*}||\leq||d||^{2}||d^{-1}||^{2}.$
Thus the family $\{x_{g}\}_{g\in G}$ satisfies (A). $\Box$
Now, let $\varphi$ be a normal faithful $G$-strongly quasi-invariant state on $\mathcal{A}$ with cocycle $x_{g}^{\prime}$s satisfying assumption (A). Since $x_{g}$ is a positive operator, one has
$x_{g}\leq\lambda,\quad\forall g\in G.$
Thanks to (2.3), this is equivalent to
$x_{g^{-1}}\geq\frac{1}{\lambda},\quad\forall g\in G.$ (3.2)
Therefore it is clear that assumption (A) is equivalent to
$\frac{1}{\lambda}\leq x_{g}\leq\lambda,\quad\forall g\in G.$ (3.3)
As a consequence of Theorem 3.2, we prove the following.
Theorem 3.3
$\varphi$ is a normal faithful $G$-strongly quasi-invariant state with cocycle $x_{g}$’s satisfying (A) if and only if there exist a normal faithful $G$-invariant state $\psi$ and a bounded positive invertible operator $d\in\mathcal{A}$ such that for all $g\in G$
$\psi(a):=\varphi(da),\quad a\in\mathcal{A}.$
Moreover, in this case for all $g\in G$, $[d,\;g(d)]=0$.
Proof. Since $\varphi$ is a normal faithful $G$–strongly quasi-invariant state with cocycle $x_{g}$’s satisfying assumption (A) then by Theorem 3.2, there exist a normal faithful $G$-invariant state $\psi$ and a bounded operator $d\in\mathcal{A}$ such that for all $g\in G$
$\psi(a):=\varphi(da),\quad a\in\mathcal{A}.$
Remember that from the proof of Theorem 3.2, $K$ is the $\sigma$-weak closure of $\mbox{conv}\{x_{g}\}_{g\in G}$ and the cocycle $x_{g}$’s satisfy (3.3), it follows that $d$ satisfies (3.3) and $d$ is a positive invertible operator in $\mathcal{A}$. Notice that in this case $x_{g}=dg^{-1}(d^{-1})=x_{g}^{*}$ which is equivalent to $[d,\;g(d)]=0$.
Conversely, let $\psi$ be a normal faithful $G$-invariant state
$\psi(a):=\varphi(da),\quad a\in\mathcal{A}$
for some normal state $\varphi$ on $\mathcal{A}$, where $d$ is a positive invertible operator in $\mathcal{A}$ such that for all $g\in G$, $[d,\;g(d)]=0$. Then it is clear that $\varphi$ is a normal faithful $G$–strongly quasi-invariant state on $\mathcal{A}$ with cocycle $x_{g}=dg^{-1}(d^{-1})=x_{g}^{*}$. Since $d$ is a positive invertible operator, it is clear that the cocycle $x_{g}^{\prime}$s satisfies (A). $\Box$
4 Unitary implementation associated to quasi–invariant states
When $\varphi$ is $G$-strongly quasi-invariant state on a von Neumann algebra $\mathcal{A}$ (for a given action) with cocycle $x_{g}$’s, it is proved in [2] that, in the GNS representation of $\varphi$, there exists a unique unitary representation of $G$ implementing the action by conjugation. We will recover it in this section and show that for $G$-quasi-invariant states, one can always spatially implement the action of $G$ (but we don’t know if the map $g\mapsto U_{g}^{*}$ is a representation). Note that in [10], it is showed that the group $G$ has always a canonical unitary implementation which is different to the one introduced here. In particular in [Theorem 3.2, [10]], with the usual notation from there, the condition $u_{g}J=Ju_{g}$ is not satisfied in our case.
Let $G$ be a group of normal *-automorphisms of a von Neumann algebra $\mathcal{A}$ and $\varphi$ be a normal $G$-quasi-invariant state on $\mathcal{A}$ with cocycle $x_{g}^{\prime}$s.
We will use Haagerup formalism in this section for the standard form of $\mathcal{A}$ as in section II of [23], this will avoid heavy formulas. We denote by $\mathcal{N}$ the crossed product $\mathcal{A}\rtimes_{\sigma^{\varphi}}\mathbb{R}$ of $\mathcal{A}$ by the modular group $\sigma^{\varphi}$ associated with $\varphi$. We denote by $\theta$ the dual action of $\mathbb{R}$ on $\mathcal{N}$. Then $\theta$ satisfies $\theta_{s}(x)=x\;\;(s\in\mathbb{R})$ iff $x\in\mathcal{A}$. Let $\tau$ be the trace on $\mathcal{N}$ such that $\tau\circ\theta_{s}=e^{-s}\tau$ and denote by $L_{0}(\mathcal{N},\;\tau)$ the space of $\tau$-measurable operators $x$ affiliated with $\mathcal{N}$. Define for $p=1,2$
$L_{p}(\mathcal{A}):=\{x\in L_{0}(\mathcal{N},\;\tau)|\;\;\theta_{s}(x)=e^{-s/p}x,\;\;s\in\mathbb{R}\}.$
There is an identification of $\mathcal{A}_{*}$ with $L_{1}(\mathcal{A})$ assigning a positive operator $D_{\gamma}\in L_{1}(\mathcal{A})$ to a normal linear form $\gamma\in\mathcal{A}_{*}$ (Proposition 15 [23]). The duality is defined through the use of a trace-like functional $\mbox{tr}:L_{1}(\mathcal{A})\to\mathbb{C}$ with
$\langle D_{\gamma},\;a\rangle_{L_{1}(\mathcal{A}),\;\mathcal{A}}=\mbox{tr}(D_{\gamma}a)=\gamma(a).$
Since $\varphi$ is faithful state on $\mathcal{A}$, its density $D_{\varphi}\in L_{1}(\mathcal{A})$ satisfies supp$(D_{\varphi})=1$ by Proposition 4(2c) in [23]. Note that, by duality, $G$ acts on $L_{1}(\mathcal{A})$ and we have
$\langle g^{*}D_{\varphi},\;a\rangle_{L_{1}(\mathcal{A}),\,\mathcal{A}}=\varphi(g(a)).$
The space $L_{2}(\mathcal{A})$ is naturally a Hilbert space with scalar product $\langle A,B\rangle=\mbox{tr}{(A^{*}B)}$. Moreover, by Theorem 36 in [23], the GNS representation of $(\mathcal{A},\;\varphi)$ is the triple $(\mathcal{A}\subset\mathcal{N},\;L_{2}(\mathcal{A}),\;D_{\varphi}^{1/2})$.
Lemma 4.1
There exists a positive invertible operator $a_{g}$ in $\mathcal{A}$ such that
$D_{\varphi\circ g}=D_{\varphi}^{1/2}a_{g}^{2}D_{\varphi}^{1/2}=x_{g}^{*}D_{\varphi}.$ (4.1)
Proof. Since $\varphi$ is normal $G$-quasi-invariant and for all $a\in\mathcal{A}$ and $g\in G$,
$\varphi(x_{g}a)=\varphi(ax_{g}^{*})$, it follows that
$g^{*}D_{\varphi}:=D_{\varphi\circ g}=x_{g}^{*}D_{\varphi}=D_{\varphi}x_{g}.$ (4.2)
On the other hand, for any positive operator $a\in\mathcal{A}$ and any $g\in G$, one has
$\varphi(g(a))=\mbox{tr}{(D_{\varphi\circ g}\,a)}=\varphi(x_{g}a)\leq||x_{g}||\varphi(a)=||x_{g}||\,\mbox{tr}{(D_{\varphi}^{1/2}aD_{\varphi}^{1/2})}.$
By Theorem 7 [23], we have $0\leq D_{\varphi\circ g}\leq\|x_{g}\|D_{\varphi}$. As $D_{\varphi}$ has full support, there exists a positive element $a_{g}$ of $\mathcal{N}$ such that
$D_{\varphi\circ\,g}=D_{\varphi}^{1/2}a_{g}^{2}D_{\varphi}^{1/2}.$
Considering the dual action yields that $\theta_{s}(a_{g})=a_{g}$ for all $s\in\mathbb{R}$ and $a_{g}$ is a positive operator in $\mathcal{A}$. Now, notice that by the proof of Lemma 3.1, $\varphi\leq||x_{g^{-1}}||\,\varphi\circ g$. Hence one gets
$D_{\varphi\circ g}=D_{\varphi}^{1/2}a_{g}^{2}D_{\varphi}^{1/2}\geq\alpha D_{\varphi},$
where $\alpha=||x_{g^{-1}}||^{-1}$. This proves that $a_{g}^{2}\geq\alpha$ and $a_{g}$ is invertible. $\Box$
Actually, the equation $x_{g}^{*}D_{\varphi}=D_{\varphi}x_{g}$ implies that $x_{g}$ sits in the domain of $\sigma^{\varphi}_{-i}$ and $\sigma^{\varphi}_{-i}(x_{g})=x_{g}^{*}$ (see Theorem 3.25 in [22]). In particular, $x_{g}$ sits in the domain of $\sigma^{\varphi}_{-i/2}$ and one can check that
$a_{g}^{2}=\sigma^{\varphi}_{-i/2}(x_{g}).$ (4.3)
Now we prove the following.
Theorem 4.2
For all $g\in G$, the map
$U_{g}:$ $\,\mathcal{A}\,D_{\varphi}^{1/2}\subset L_{2}(\mathcal{A})\longrightarrow L_{2}(\mathcal{A})$
$xD_{\varphi}^{1/2}\longmapsto g^{-1}(x)D_{\varphi}^{1/2}a_{g}$
extends to a unitary transformation on $L_{2}(\mathcal{A})$ (still denoted by $U_{g}$). Moreover for any $x\in\mathcal{A}$
$U_{g}^{*}xU_{g}=g(x).$
Proof. For all $x,y\in\mathcal{A}$, one has
$\langle U_{g}(xD_{\varphi}^{1/2}),\,U_{g}(yD_{\varphi}^{1/2})\rangle_{L_{2}(\mathcal{A})}$ $=$ $\mbox{tr}{(a_{g}D_{\varphi}^{1/2}g^{-1}(x)^{*}g^{-1}(y)D_{\varphi}^{1/2}a_{g})}$
$=$ $\mbox{tr}{(D_{\varphi}^{1/2}a_{g}^{2}D_{\varphi}^{1/2}g^{-1}(x)^{*}g^{-1}(y))}$
$=$ $\varphi(x^{*}y)$
$=$ $\langle xD_{\varphi}^{1/2},\,yD_{\varphi}^{1/2}\rangle_{L_{2}(\mathcal{A})}.$
Since $\mathcal{A}\,D_{\varphi}^{1/2}$ is dense in $L_{2}(\mathcal{A})$, $U_{g}$ can be extended to an isometry on $L_{2}(\mathcal{A})$. Now, let $h\in L_{2}(\mathcal{A})\cap\mbox{Ran}(U_{g})^{\bot}$. Then, for all $a\in\mathcal{A}$
$0=\langle h,\,U_{g}(aD_{\varphi}^{1/2})\rangle_{L_{2}(\mathcal{A})}=\mbox{tr}{(h^{*}g^{-1}(a)D_{\varphi}^{1/2}a_{g})}=\mbox{tr}{(a_{g}h^{*}g^{-1}(a)D_{\varphi}^{1/2})}.$
Notice that $\mathcal{A}\,D_{\varphi}^{1/2}$ is dense in $L_{2}(\mathcal{A})$ and $a_{g}$ is invertible. Therefore, one gets $a_{g}h^{*}=0$ and $h=0$. Hence, $U_{g}$ is a unitary operator on $L_{2}(\mathcal{A})$.
Now, let $x,y,z\in\mathcal{A}$. Then, one has
$\langle U_{g}^{*}xU_{g}(yD_{\varphi}^{1/2}),\,zD_{\varphi}^{1/2}\rangle_{L_{2}(\mathcal{A})}$ $=$ $\langle xg^{-1}(y)D_{\varphi}^{1/2}a_{g},\;g^{-1}(z)D_{\varphi}^{1/2}a_{g}\rangle_{L_{2}(\mathcal{A})}$
$=$ $\mbox{tr}{(a_{g}D_{\varphi}^{1/2}(g^{-1}(y))^{*}x^{*}g^{-1}(z)D_{\varphi}^{1/2}a_{g})}$
$=$ $\varphi(y^{*}g(x)^{*}z)$
$=$ $\langle g(x)yD_{\varphi}^{1/2},\;zD_{\varphi}^{1/2}\rangle_{L_{2}(\mathcal{A})},$
which proves that $U_{g}^{*}xU_{g}=g(x).$Box
In the setting of Theorem 3.2, we have that there exists an invertible $a\in\mathcal{A}$ such that $D_{\psi}=D_{\varphi}^{1/2}a^{2}D_{\varphi}^{1/2}$. Thus there is a partial isometry $v$ such that $D_{\psi}^{1/2}=vaD_{\varphi}^{1/2}$. Since both operators have full support, $v$ is unitary and we deduce that there exists an invertible $\gamma\in\mathcal{A}$ such that $D_{\psi}^{1/2}=\gamma D_{\varphi}^{1/2}$. Thus we get $D_{\psi}=d^{*}D_{\varphi}=\gamma D_{\varphi}\gamma^{*}$. As above, we get that $\gamma^{*}$ sits in the domain of $\sigma^{\varphi}_{-i}$ and $\sigma^{\varphi}_{-i}(\gamma^{*})=\gamma^{-1}d^{*}$. We conclude that $d^{*}=\gamma\sigma^{\varphi}_{-i}(\gamma^{*})$. In the same way, let $\gamma_{g}=g^{-1}(\gamma^{-1})\gamma$ for $g\in G$, then one can also check that $\gamma_{g}^{*}$ sits in the domain of $\sigma^{\varphi}_{-i}$ and that $x_{g}^{*}=\gamma_{g}\sigma^{\varphi}_{-i}(\gamma_{g}^{*})$.
Proposition 4.3
If $\varphi$ is $G$-strongly quasi-invariant then $g\mapsto U_{g}^{*}$ is a unitary representation of the group $G$.
Proof. Notice that if $\varphi$ is $G$-strongly quasi-invariant, it is showed in [2] that the cocycle $x_{g},\;g\in G$, are in the centralizer of $\varphi$. Therefore by (4.3) it follows that $a^{2}_{g}=x_{g}=x_{g}^{*}$ (i.e. $a_{g}=x_{g}^{1/2}$). Moreover from (4.2), one has
$a_{g}D_{\varphi}^{1/2}=D_{\varphi}^{1/2}a_{g}$
and hence for all $x\in\mathcal{A}$, $g\in G$
$U_{g}(xD_{\varphi}^{1/2})=g^{-1}(x)x_{g}^{1/2}D_{\varphi}^{1/2}.$
By using the cocycle property, for any $g,h\in G$, $x\in\mathcal{A}$, one has
$U_{g}U_{h}(xD_{\varphi}^{1/2})=U_{g}(h^{-1}(x)x_{h}^{1/2}D_{\varphi}^{1/2})=g^{-1}(h^{-1}(x))g^{-1}(x_{h}^{1/2})x_{g}^{1/2}D_{\varphi}^{1/2}=U_{hg}(xD_{\varphi}^{1/2}).$
It follows that $g\mapsto U_{g}^{*}=U_{g^{-1}}$ is a unitary representation of the group $G$. $\Box$
5 A Theorem of Kovács and Szücs for quasi–invariant states with uniformly bounded cocycle
By Theorem 3.2, if $\varphi$ is a normal faithful $G$-quasi–invariant state on a von Neumann algebra $\mathcal{A}$ with cocycle satisfying assumption (A), then there exist a normal faithful $G$-invariant state $\psi$ on $\mathcal{A}$ and a bounded operator $d\in\mathcal{A}$ such that for all $g\in G$
$\psi(a):=\varphi(da),\quad a\in\mathcal{A}.$
Since $\psi$ is a normal faithful $G$-invariant state on $\mathcal{A}$, it follows from [8, Theorem 1] that there exists a normal faithful $G$-invariant Umegaki conditional expectation $\Phi:\mathcal{A}\rightarrow\mathcal{B}$, where $\mathcal{B}$ is a sub-von Neumann algebra of $\mathcal{A}$, such that
$\psi(.)=(\psi|\mathcal{B})\circ\Phi(.)$
If moreover $d$ is invertible, one gets
$\varphi=(\varphi(d.)|\mathcal{B})\circ\Phi(d^{-1}.)$ (5.1)
Now, if $\varphi$ is a $G$-strongly quasi-invariant state and if $g\mapsto U_{g}^{*}$ is the unitary representation of $G$ defined by Theorem 4.2, we will show that the conditional expectation $\Phi$ in (5.1) can also be characterized by $U_{g},\;g\in G$ Define
$u_{g}(a):=U^{*}_{g}aU_{g}=g(a)\ ;\quad\forall\,g\in G\ ;\quad\forall\,a\in\mathcal{A}$ (5.2)
Let $E_{0}$ be the orthogonal projection onto the subspace of all $\xi\in L_{2}(\mathcal{A})$ invariant under all $U_{g}$. We denote by $\mathcal{B}$ the set of fixed points of $\mathcal{A}$ under the action of the automorphisms $u_{g}$, $g\in G$. Let $F_{0}=[\mathcal{B}^{\prime}E_{0}]$ be the orthogonal projection onto $\overline{\mbox{span}(\mathcal{B}^{\prime}E_{0}L_{2}(\mathcal{A}))}$. Finally, let $\mathcal{M}$ be a Godement mean over $G$. The following result follows from a theorem of Kovács and Szücs (see [14] or [1] and [8, Theorem 1]).
Theorem 5.1
Assume that $\varphi$ is a faithful normal $G$-strongly quasi-invariant state with uniformly bounded cocycle (i.e. cocycle satisfying (A)). Then
  1. (1)
    There exists a unique normal faithful $G$-invariant conditional expectation
    $\Phi:\mathcal{A}\rightarrow\mathcal{B}$. Moreover, $\Phi(b)$ can be defined (equivalently) as
    1. (i)
      the unique element of $\mathcal{A}$ such that $\Phi(b)E_{0}=E_{0}bE_{0}$
    2. (ii)
      the unique element of $\mathcal{A}$ such that
      $(\xi_{1},\;\Phi(b)\xi_{2})=\mathcal{M}\{(\xi_{1},\;U_{g^{-1}}bU_{g}\xi_{2})\}$
  2. (2)
    A faithful normal state $w$ on $\mathcal{A}$ is $G$-strongly quasi-invariant with uniformly bounded cocycle if and only if there exists a positive invertible bounded operator $d$ such that $w(.)=(w(d.)|\mathcal{B})\circ\Phi(d^{-1}.)$
Proof. Let $\varphi$ is a faithful normal $G$-strongly quasi-invariant state with uniformly bounded cocycle. By applying Theorem 3.3, there exists a bounded positive invertible operator $d$ and a state a $G$-invariant state $\psi$ such that
$\psi(a)=\varphi(da)=\langle d^{1/2}D_{\varphi}^{1/2},ad^{1/2}D_{\varphi}^{1/2}\rangle_{L_{2}(\mathcal{A})}\quad\forall a\in\mathcal{A}.$
Remember that $x_{g}=dg^{-1}(d^{-1})=x_{g}^{*}$. Moreover for all $g\in G$
$U_{g}d^{1/2}D_{\varphi}^{1/2}=g^{-1}(d^{1/2})x_{g}^{1/2}D_{\varphi}^{1/2}=d^{1/2}D_{\varphi}^{1/2}.$
Let $p$ be the orthogonal projection onto $\overline{\mbox{span}(\mathcal{B}^{\prime}d^{1/2}D_{\varphi}^{1/2}})$. Then $p\in\mathcal{B}$ and $p\leq F_{0}$. Moreover, one has
$\psi(p)$ $=$ $\langle d^{1/2}D_{\varphi}^{1/2},pd^{1/2}D_{\varphi}^{1/2}\rangle_{L_{2}(\mathcal{A})}$
$=$ $\langle d^{1/2}D_{\varphi}^{1/2},d^{1/2}D_{\varphi}^{1/2}\rangle_{L_{2}(\mathcal{A})}$
$=$ $\varphi(d)=\psi(I)=1.$
Since $\psi$ is faithful, $p=I$ and $F_{0}=I$. Finally the results of the above theorem follows from [8, Theorem 1]. $\Box$
Now we provide a commutative easy example of faithful normal $G$-quasi-invariant state such that the cocycle is not uniformly bounded. Moreover, we show in this example that there is no $G$-invariant conditional expectation.
Example. On $\mathcal{A}=L^{\infty}(\mathbb{R})$, define the state
[…]
The additive group $\mathbb{R}$ acts on $\mathcal{A}$ by translation:
$\tau_{t}(f)(s):=f(s-t)$
It is clear that $\varphi$ is $\mathbb{R}$-quasi-invariant state with bounded cocycle
$x_{t}(s)=\frac{1+s^{2}}{1+(s+t)^{2}}.$
The cocycle $(x_{t})_{t}$ does not satisfy assumption (A) since $\|x_{t}\|,\|x_{t}^{-1}\|\geq 1+t^{2}$. The fixed point of the action are constant functions. Thus every conditional expectation $\Phi$ on $L^{\infty}(\mathbb{R})$ has to be of the form
$\Phi(f)=\psi(f)1$ (5.3)
for some normal state $\psi$. But there are no functions in $L_{1}$ that are $\mathbb{R}$-invariant.
This example also shows that it seems unlikely to improve Theorem 3.2.
6 Strongly quasi-invariant states and semifinite trace
In this section we extend the result developped by $\o$rmer [20] to the case of strongly quasi-invariant states with uniformly bounded cocycles. Given a semifinite von Neumann algebra $\mathcal{A}$, a group $G$ of *-automorphisms is said to act ergodically on the center $\mathcal{Z}=\mathcal{A}\cap\mathcal{A}^{\prime}$ of $\mathcal{A}$ if $\mathcal{F}(G)\cap\mathcal{Z}=\mathbb{C}1$, where $\mathcal{F}(G)$ is the set of the fixed points of $G$ in $\mathcal{A}$.
Theorem 6.1
Let $\mathcal{A}$ be a semifinite von Neumann algebra and let $G$ be an amenable group of *-automorphisms of $\mathcal{A}$ and assume that $G$ acts ergodically on the center $\mathcal{Z}$ of $\mathcal{A}$. Suppose that $\varphi$ is a faithful normal $G$-quasi-invariant state on $\mathcal{A}$ with cocycle $x_{g}$’s satisfying assumption $(A)$. Then, there exists up to a scalar multiple a unique faithful normal $G$-invariant semifinite trace $\tau$ of $\mathcal{A}$. Furthermore the density $c\in L_{1}(\mathcal{A},\tau)^{+}$ of $\varphi$ with respect to $\tau$ satisfies
$(g^{-1})^{*}(c)=cx_{g^{-1}},\quad x_{g}^{*}c=cx_{g},\quad\forall g\in G,$ (6.1)
where $(g^{-1})^{*}:L_{1}(\mathcal{A},\tau)\rightarrow L_{1}(\mathcal{A},\tau)$ is the predual map of $g^{-1}$ on $\mathcal{A}$.
Proof. Under assumptions of the above theorem, it follows from Theorem 3.2 that there exists a normal faithful $G$-invariant state $\psi$. Applying [[20], Theorem 1] for $\psi$, there exists up to a scalar multiple a unique faithful normal $G$-invariant semifinite trace $\tau$ on $\mathcal{A}$. Let $c\in L_{1}(\mathcal{A},\tau)^{+}$ be the density of $\varphi$. For all $a\in\mathcal{A}$, we have
$\tau((g^{-1})^{*}(c)a)=\tau(cg^{-1}(a))=\varphi(g^{-1}(a))=\varphi(x_{g^{-1}}a)=\tau(cx_{g^{-1}}a).$
This proves the first equality (6.1). For the second, note that for all $a\in\mathcal{A}$
$\tau(cx_{g}a)=\varphi\circ g(a)=\overline{\varphi\circ g(a^{*})}=\overline{\tau(cx_{g}a^{*})}=\tau(x_{g}^{*}ca).$
This implies that $x_{g}^{*}c=cx_{g}$ in $L_{1}(\mathcal{A},\tau)$. $\Box$
Example. Taking again $\mathcal{A}=L^{\infty}(\mathbb{R})$ and $G$ to be the $ax+b$ group of all automorphisms of $\mathcal{A}$ of the form $\pi_{(a,b)}f(t)=f(at+b)$ for $(a,b)\in\mathbb{R}^{*}\times\mathbb{R}$. Then $G$ is amenable (since it is solvable), it admits a cocycle for the state given by the measure $\frac{1}{\pi}\frac{1}{1+t^{2}}dt$; $x_{(a,b)}(t)=\frac{a(1+t^{2})}{a^{2}+(t-b)^{2}}$ that is not uniformly bounded. Moreover, one can check that the fixed point algebra by $G$ consists only in constant functions. Similarly, one also easily sees that there are no measures on $\mathbb{R}$ invariant by $G$. Thus one cannot remove the assumption $(A)$ even in a commutative situation.
Acknowledgement
A. Dhahri is a member of GNAMPA-INdAM and he has been supported by the MUR grant Dipartimento di Eccellenza 2023-2027 of Dipartimento di Matematica, Politecnico di Milano.
Statements and declarations
Ethical Approval: Not applicable.
Competing interests: The authors declare that they have no conflict of interest.
Data availability statement: The manuscript has no associated data.